OpenStudy (anonymous):

I need help arranging steps to an equation: Vx+3 - V2x-1=2 The Vs represent square root symbols. These are my options. 1.)Simplify to obtain the final radical term on one side of the equation. 2.)Raise both sides of the equation to the power of 2. 3.)Apply the Zero Product Rule. 4.)Use the quadratic formula to find the values of x. 5.)Simplify to get a quadratic equation. 6.)Raise both sides of the equation to the power of 2 again. I'd be grateful if someone could give me a short tutoring on how this exactly works.

OpenStudy (michele_laino):

I think that we can rewrite your equation as follows: \[\large \sqrt {x + 3} = 2 + \sqrt {2x - 1} \]

OpenStudy (anonymous):

Alright, I'm writing this down.

OpenStudy (anonymous):

Raising the sides by 2 right? That's the first step?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

step#2 we have to square both sides, so we get: \[\large x + 3 = 4 + 2x - 1 + 4\sqrt {2x - 1} \]

OpenStudy (anonymous):

Good, and then from the looks of it we have to simplify.

OpenStudy (anonymous):

Oh, you already put it down.

OpenStudy (michele_laino):

step#3 we have to simplify that equation, as follows:

OpenStudy (anonymous):

Into a quadratic equation yes? It's starting to come back to me.

OpenStudy (michele_laino):

\[\large \begin{gathered} x + 3 - 4 - 2x + 1 = 4\sqrt {2x - 1} \hfill \\ - x = 4\sqrt {2x - 1} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

step#4 we have to square both sides again:

OpenStudy (michele_laino):

\[\large {x^2} = 16\left( {2x - 1} \right)\]

OpenStudy (anonymous):

After that there should be one final step.

OpenStudy (michele_laino):

step#5 I simplify that equation as follows: \[\large {x^2} - 32x + 16 = 0\]

OpenStudy (anonymous):

And that's to simplify.

OpenStudy (michele_laino):

yes! We can write:

OpenStudy (michele_laino):

step#6 \[\large \begin{gathered} {x_1} = 16 + 4\sqrt {15} \hfill \\ {x_2} = 16 - 4\sqrt {15} \hfill \\ \end{gathered} \] those are the solutions of the last equation

OpenStudy (anonymous):

So the Zero Product rule is no concern here?

OpenStudy (anonymous):

Or does that go with step 4?

OpenStudy (michele_laino):

namely I have used the quadratic formula

OpenStudy (anonymous):

Ah, now I understand it.

OpenStudy (anonymous):

I think I'm capable of starting my Algebra Test. Wish me luck.

OpenStudy (michele_laino):

we have to verify if those roots, are really the solutions of your irrational equation

OpenStudy (michele_laino):

those numbers are solutions to your irrational equation, if they satisfy the subsequent inequalities: \[\large \begin{gathered} x + 3 \geqslant 0 \hfill \\ 2x - 1 \geqslant 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I passed!

OpenStudy (michele_laino):

I can assure that both those numbers verify those inequalities, so the solutions of your irrational equation, are: \[\large \begin{gathered} {x_1} = 16 + 4\sqrt {15} \hfill \\ {x_2} = 16 - 4\sqrt {15} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

do you need more help?

OpenStudy (anonymous):

Possibly in the future, as far as I'm concerned. For now I'm off to study on Polynomial Functions.

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thank you so much! I hope you have a great day!

OpenStudy (michele_laino):

Thank you! :)